A=yz then it's gradient is






1
Expert's answer
2022-01-10T09:17:00-0500

The gradient of a scalar field is given by

"{\rm grad} A=\hat i \frac{\partial A}{\partial x}+\hat j \frac{\partial A}{\partial y}+\hat k \frac{\partial A}{\partial z}"

In our case

"{\rm grad} A=\hat i \frac{\partial (yz)}{\partial x}+\hat j \frac{\partial (yz)}{\partial y}+\hat k \frac{\partial (yz)}{\partial z}"

"{\rm grad} A=\hat j z+\hat k y"